# Shadows of the Colossus: Hierarchical Black Hole Mergers in a 10-million-body Globular Cluster Simulation

### v2

@@ 4.3 Ejected BBH properties
This trend also explains why the majority (roughly 80%) of the BBH mergers in colossus occur inside the cluster (see Table 1). The GW inspiral time of a BBH scales steeply with semi-major axis ($t_{\rm insp}\propto a^{4}$; Peters1964). This means BBHs in more massive clusters that are near the critical orbital separation for dynamical ejection are more susceptible to merging inside their host before ejection can occur (see also, e.g., Rodriguez2019; Kremer2020; AntoniniGieles2020). Indeed, this explains the trend shown in the right-panel of Figure 7 where the ratio of in-cluster to ejected mergers increases with cluster mass.
5 Mapping to Virgo globular clusters
With the addition of colossus, the CMC Cluster Catalog now adequately spans the mass-$r_{h}$ parameter space of observed GCs in the Virgo Supercluster. This enables extrapolation of quantities like the total number of binary BH mergers and BBH merger rate from the CMC catalog to observed Virgo clusters, and prediction of total BBH merger rate from GC sources in the Virgo supercluster as a whole. We introduce a mass-$r_{h}$ distance function to map Virgo GCs to their closest CMC catalog counterparts, and utilize the scaling relations in cluster mass and $r_{h}$ derived in Section 4 to predict BBH merger quantities for each Virgo cluster, based on its closest CMC counterparts. Unless otherwise specified, all CMC model properties referred to in this section are present-day values ($t>12\,\text{Gyr}$).
5.1 Cluster mapping with distance function
@@
For Virgo GC properties, we use the results from the “ACS Virgo Cluster Survey” (e.g., Cote2004). In particular, we use the catalog in Jordan2009 to attain present-day cluster masses and half-light radii.
Figure 11: Estimated total BBH merger populations ($N_{\rm bbh}$, left panel) and generation 2+ merger populations ($N_{\rm 2G+}$, right panel) for each GC in the Virgo cluster, utilizing the scaling relations and weighting scheme described in the text. Constant curves for 1, 10, 100, and 1000 mergers are shown in both panels, obtained by scaling off of a middling CMC model with $\log(M_{\rm cluster}/M_{\odot})=5.7$, $r_{h}=5.0\,\text{pc}$, $N_{\rm bbh}=153$, and $N_{\rm 2G+}=14$ (model 104 in kremer2020modeling). As expected, merger counts generally grow with increasing cluster mass and decreasing $r_{h}$.
@@
5.2 Interpolating the CMC model grid
Using Equation 7 to quantify the similarity between CMC models and Virgo clusters, we can utilize the scaling relations of Section 4 to interpolate the results of the CMC models and predict the number of mergers for each observed Virgo GC, and also extrapolate to Virgo GCs outside of the region covered by the CMC model parameter space.
In Section 4, we derived the scaling relations $N_{\rm bbh}~\propto~M_{\rm cl}^{1.4}$ (Equation 3) and $N_{\rm bbh}~\propto~e^{-0.2r_{h}}$ (Equation 5) in cluster mass $M$ and half-light radius $r_{h}$. Since the mass relation is independent of half-light radius and vice-versa, we combine the two relations as a separable function of the form
@@ 5.2 Interpolating the CMC model grid
We can also estimate the number of generation 2+ hierarchical mergers in a Virgo cluster, $N_{\rm 2G+}$, utilizing the $N_{\rm 2G+}\propto M^{2}$ and $N_{\rm 2G+}\propto e^{-0.4r_{h}}$ relations from Section 4 to modify Equation 9. This amounts to simply replacing the mass power with $2$ and the $r_{h}$ decay constant with 0.4-0.4, as well as $N_{\rm bbh}^{\texttt{CMC}}$ with $N_{\rm 2G+}^{\texttt{CMC}}$.
@@
Using the weighting scheme of Equation 10, a CMC model that is twice as far from the target Virgo GC as the closest model will be weighted half as much, thrice as far will be weighted one-third as much, etc. An equivalent method can be used to calculate $N_{\rm 2G+}^{\rm Virgo}$, simply by replacing $N_{n}^{\rm bbh}$ with $N_{n}^{\rm 2G+}$ (obtained using the appropriate 2G+ scaling relations). Since Equation 9 may break down when extrapolating over vast mass or $r_{h}$ ranges, only the five closest CMC models are considered, and Virgo GCs which exceed the initial mass of colossus are excluded from the rest of the analysis ($\log M_{\rm cl}>6.8$, only six GCs in the Jordan2009 catalog).
5.3 Predicted Virgo cluster BBH merger populations
We utilize Equation 10 to calculate $N_{\rm bbh}$ and $N_{\rm 2G+}$ for each GC in the Virgo cluster. Figure 11 again plots the mass-$r_{h}$ parameter space of Virgo GCs (excluding six GCs with mass exceeding the initial mass of colossus), colored according to the predicted number of total BBH mergers ($N_{\rm bbh}$, left panel) and the predicted number of generation 2+ hierarchical mergers ($N_{\rm 2G+}$, right panel). The total number of mergers in each panel is shown to increase as the cluster density increases, as expected. For very low-mass clusters with present-day masses less than roughly $10^{4.5}\,M_{\odot}$, the predicted number of mergers can be seen to increase slightly. This is because some of these clusters are best matched by CMC models that were initially more massive (and therefore formed many BBH mergers, especially at early times), but lost a significant amount of mass via enhanced tidal stripping in their galactic potential (see kremer2020modeling, for further discussion).
@@ 5.3 Predicted Virgo cluster BBH merger populations
6 Summary and conclusions
6.1 Summary
@@
The output from colossus is now publicly available as part of the CMC Cluster Catalog. [^4] https://cmc.ciera.northwestern.edu Additionally, the Virgo interpolation scheme developed in this study is available on GitHub, [^5] https://github.com/aidanmai/colossus along with all BBH merger data from colossus, and notebooks for generating the figures in Section 5.
6.2 Future Work
@@
We have focused here on black hole growth via hierarchical mergers, however other growth channels are likely possible in dense stellar clusters. One scenario is growth via massive stellar mergers and collisions while the cluster is very young ($t\lesssim 10\,$Myr). A number of studies have demonstrated these may provide a pathway for forming black holes in or beyond the pair-instability mass gap (e.g., DiCarlo2019; Kremer2020; Gonzalez2021; Costa2022; Ballone2023). Indeed, $767$ black holes with masses in the pair-instability gap formed via binary stellar collisions/mergers in colossus (see the small handful of G1 points in Figure 3 with primary mass $40.5\,M_{\odot}$ which formed via this channel). Additionally, several studies have shown that black holes may also grow via accretion upon collision with other massive stars (e.g., Giersz2015; Kiroglu2025a; Kiroglu2025c). Rose2023 pointed out that this channel may be especially prominent in very dense star clusters, with densities similar to the Galactic center. In colossus, we identify $759$ black hole+star collisions events over the cluster’s lifetime. For the vast majority of these ($90\%$), the $M_{\star}/M_{\rm bh}$ mass ratio is $0.1$ or less. In this regime, significant black hole growth is unlikely (Kremer2022_sph). Only five collisions feature $M_{\star}/M_{\rm bh}$ values of 0.5 or higher (all occurring before 2525\,Myr) where significant growth may be feasible. However, we have assumed in colossus a $5\%$ binary fraction for all stars. Realistic clusters may very well have much higher binary fractions, especially for massive stars (e.g., Offner2023). Higher binary fractions may significantly increase the number of massive star–black hole collisions (Kiroglu2025a). We reserve exploration of this possibility for future models.

### v3

@@ IV.3 Ejected BBH properties
This trend also explains why the majority (roughly 80%) of the BBH mergers in colossus occur inside the cluster (see Table 1). The GW inspiral time of a BBH scales steeply with semi-major axis ($t_{\rm insp}\propto a^{4}$; Peters, 1964). This means BBHs in more massive clusters that are near the critical orbital separation for dynamical ejection are more susceptible to merging inside their host before ejection can occur (see also, e.g., Rodriguez et al., 2019; Kremer and others, 2020; Antonini and Gieles, 2020). Indeed, this explains the trend shown in the right-panel of Figure 7 where the ratio of in-cluster to ejected mergers increases with cluster mass.
V Mapping to Virgo globular clusters
With the addition of colossus, the CMC Cluster Catalog now adequately spans the mass-$r_{h}$ parameter space of observed GCs in the Virgo Supercluster. This enables extrapolation of quantities like the total number of binary BH mergers and BBH merger rate from the CMC catalog to observed Virgo clusters, and prediction of total BBH merger rate from GC sources in the Virgo supercluster as a whole. We introduce a mass-$r_{h}$ distance function to map Virgo GCs to their closest CMC catalog counterparts, and utilize the scaling relations in cluster mass and $r_{h}$ derived in Section IV to predict BBH merger quantities for each Virgo cluster, based on its closest CMC counterparts. Unless otherwise specified, all CMC model properties referred to in this section are present-day values ($t>12\,\text{Gyr}$).
V.1 Cluster mapping with distance function
@@
For Virgo GC properties, we use the results from the “ACS Virgo Cluster Survey” (e.g., Côté et al., 2004). In particular, we use the catalog in Jordán et al. (2009) to attain present-day cluster masses and half-light radii.
Figure 11: Estimated total BBH merger populations ($N_{\rm bbh}$, left panel) and generation 2+ merger populations ($N_{\rm 2G+}$, right panel) for each GC in the Virgo cluster, utilizing the scaling relations and weighting scheme described in the text. Constant curves for 1, 10, 100, and 1000 mergers are shown in both panels, obtained by scaling off of a middling CMC model with $\log(M_{\rm cluster}/M_{\odot})=5.7$, $r_{h}=5.0\,\text{pc}$, $N_{\rm bbh}=153$, and $N_{\rm 2G+}=14$ (model 104 in Kremer et al., 2020b). As expected, merger counts generally grow with increasing cluster mass and decreasing $r_{h}$.
@@
V.2 Interpolating the CMC model grid
Using Equation 7 to quantify the similarity between CMC models and Virgo clusters, we can utilize the scaling relations of Section IV to interpolate the results of the CMC models and predict the number of mergers for each observed Virgo GC, and also extrapolate to Virgo GCs outside of the region covered by the CMC model parameter space.
In Section IV, we derived the scaling relations $N_{\rm bbh}~\propto~M_{\rm cl}^{1.4}$ (Equation 3) and $N_{\rm bbh}~\propto~e^{-0.2r_{h}}$ (Equation 5) in cluster mass $M$ and half-light radius $r_{h}$. Since the mass relation is independent of half-light radius and vice-versa, we combine the two relations as a separable function of the form
@@ V.2 Interpolating the CMC model grid
We can also estimate the number of generation 2+ hierarchical mergers in a Virgo cluster, $N_{\rm 2G+}$, utilizing the $N_{\rm 2G+}\propto M^{2}$ and $N_{\rm 2G+}\propto e^{-0.4r_{h}}$ relations from Section IV to modify Equation 9. This amounts to simply replacing the mass power with $2$ and the $r_{h}$ decay constant with 0.4-0.4, as well as $N_{\rm bbh}^{\texttt{CMC}}$ with $N_{\rm 2G+}^{\texttt{CMC}}$.
@@
Using the weighting scheme of Equation 10, a CMC model that is twice as far from the target Virgo GC as the closest model will be weighted half as much, thrice as far will be weighted one-third as much, etc. An equivalent method can be used to calculate $N_{\rm 2G+}^{\rm Virgo}$, simply by replacing $N_{n}^{\rm bbh}$ with $N_{n}^{\rm 2G+}$ (obtained using the appropriate 2G+ scaling relations). Since Equation 9 may break down when extrapolating over vast mass or $r_{h}$ ranges, only the five closest CMC models are considered, and Virgo GCs which exceed the initial mass of colossus are excluded from the rest of the analysis ($\log M_{\rm cl}>6.8$, only six GCs in the Jordán et al. (2009) catalog).
V.3 Predicted Virgo cluster BBH merger populations
We utilize Equation 10 to calculate $N_{\rm bbh}$ and $N_{\rm 2G+}$ for each GC in the Virgo cluster. Figure 11 again plots the mass-$r_{h}$ parameter space of Virgo GCs (excluding six GCs with mass exceeding the initial mass of colossus), colored according to the predicted number of total BBH mergers ($N_{\rm bbh}$, left panel) and the predicted number of generation 2+ hierarchical mergers ($N_{\rm 2G+}$, right panel). The total number of mergers in each panel is shown to increase as the cluster density increases, as expected. For very low-mass clusters with present-day masses less than roughly $10^{4.5}\,M_{\odot}$, the predicted number of mergers can be seen to increase slightly. This is because some of these clusters are best matched by CMC models that were initially more massive (and therefore formed many BBH mergers, especially at early times), but lost a significant amount of mass via enhanced tidal stripping in their galactic potential (see Kremer et al., 2020b, for further discussion).
@@ V.3 Predicted Virgo cluster BBH merger populations
VI Summary and conclusions
VI.1 Summary
@@
The output from colossus is now publicly available as part of the CMC Cluster Catalog. [^4] https://cmc.ciera.northwestern.edu Additionally, the Virgo interpolation scheme developed in this study is available on GitHub, [^5] https://github.com/aidanmai/colossus along with all BBH merger data from colossus, and notebooks for generating the figures in Section V.
VI.2 Future Work
@@
We have focused here on black hole growth via hierarchical mergers, however other growth channels are likely possible in dense stellar clusters. One scenario is growth via massive stellar mergers and collisions while the cluster is very young ($t\lesssim 10\,$Myr). A number of studies have demonstrated these may provide a pathway for forming black holes in or beyond the pair-instability mass gap (e.g., Di Carlo and others, 2019; Kremer and others, 2020; González and others, 2021; Costa et al., 2022; Ballone et al., 2023). Indeed, $767$ black holes with masses in the pair-instability gap formed via binary stellar collisions/mergers in colossus (see the small handful of G1 points in Figure 3 with primary mass $40.5\,M_{\odot}$ which formed via this channel). Additionally, several studies have shown that black holes may also grow via accretion upon collision with other massive stars (e.g., Giersz et al., 2015; Kıroğlu et al., 2025a, b). Rose et al. (2023) pointed out that this channel may be especially prominent in very dense star clusters, with densities similar to the Galactic center. In colossus, we identify $759$ black hole+star collisions events over the cluster’s lifetime. For the vast majority of these ($90\%$), the $M_{\star}/M_{\rm bh}$ mass ratio is $0.1$ or less. In this regime, significant black hole growth is unlikely (Kremer et al., 2022). Only five collisions feature $M_{\star}/M_{\rm bh}$ values of 0.5 or higher (all occurring before 2525\,Myr) where significant growth may be feasible. However, we have assumed in colossus a $5\%$ binary fraction for all stars. Realistic clusters may very well have much higher binary fractions, especially for massive stars (e.g., Offner et al., 2023). Higher binary fractions may significantly increase the number of massive star–black hole collisions (Kıroğlu et al., 2025a). We reserve exploration of this possibility for future models.